OpenStudy (kianna0098):

Joey has to solve the system of equations using substitution. He has completed the work shown. What should be his next step?

OpenStudy (kianna0098):

A: is crossed out B: He needs to write his solution as an ordered pair (0, 2). C: He needs to plug back into one of the equations to find x. D: He needs to check his work to make sure he made no mistakes.

OpenStudy (kianna0098):

@grimesssz

OpenStudy (grimesssz):

he needs to plug back into one of the equations to find x so thus its C

OpenStudy (whpalmer4):

Well, maybe if you showed us the work shown, we could intelligently answer the question... Or, maybe an example will give you enough information to figure it out on your own. \[y = x + 2\]\[3y = 4x\]We can substitute \((x+2)\) wherever we see \(y\) in the second equation, so let's do that: \[3(x+2) = 4x\]\[3x+6 = 4x\]\[3x - 3x + 6 = 4x-3x \]\[6 = x\] Now we plug that back into our substitution equation to find \(y\): \[y = x+2 = 6 + 2 = 8\] So our answer is \((6,8)\). Now we plug our solution into both equations and verify that it satisfies both of them: \[8 = 6+2\checkmark\]\[3(8) = 4(6)\checkmark\] Our answer checks out, so \((6,8)\) is the solution to that system of equations.